# IrokoBench: A New Benchmark for African Languages in the Age of Large Language Models

### v1

@@ 4.2 Few-shot results, Cross-Lingual Transfer, Sensitivity to Prompt Templates
Figure 2: Few-shot evaluation on IrokoBench datasets, we performed 8-shots for AfriMGSM and 5-shots for the others.
@@ 5 Conclusion
Our benchmark has a few limitations: (1) The benchmark is human-translated which may include some translationese effects, it would have been better if they are all generated in the native African languages. However, this parallel translation allows us to evaluate and compare the same sentences in all these languages. (2) We only cover three language families in Africa, Nilo-Saharan, Austronesian, and Khoisan language groups are missing, one of the reason we excluded them is either lack of contact with professional translators or limited translation budget, we hope to extend to more languages in the future. (3) Prompts used for evaluation are not exhaustive some of the results may change with appropriate a more careful prompt search. The best prompt for different LLMs might be different.
@@ Acknowledgment
This work was carried out with support from Lacuna Fund, an initiative co-founded by The Rockefeller Foundation, Google.org, and Canada’s International Development Research Centre. David Adelani acknowledges the support of DeepMind Academic Fellowship programme. This work was supported in part by compute credits from Oracle and Cohere For AI Research Grant. Also, we are grateful to OpenAI for providing API credits through their Researcher Access API programme to Masakhane for the evaluation of GPT-3.5 and GPT-4 LLMs. Finally, we are grateful to Hailey Schoelkopf and Lintang Sutawika for their assistance with the EleutherAI lm-eval tool.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 5
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A]
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g., for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See Section 3.5, more details in A.2
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Section 3.5
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [N/A]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [N/A]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] See Section 3.2
    2. (b)
      Did you mention the license of the assets? [Yes] See Section 3.2
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [N/A]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] See Section 3.3 for data and 3.4 for models
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes] See Section 3.3
Appendix A Technical Appendix
@@ A.1 Language covered
Table 7 provide the languages covered in the IrokoBench, their language family, regions located in Africa, number of speakers, and size of monolingual data available on the web based on the MADLAD cleaned corpus [36]—we only report number of characters in mega bytes. Additionally, we added an indication whether this language is covered in the pre-training of Aya-101 and BLOOMZ 7B LLMs.
A.2 Evaluation Tool
@@
We use the EleutherAI LM Evaluation Harness (lm-eval) tool [15]—a popular evaluation tool that is helping to standardize LLM evaluation. The tool allows for three types of evaluation: log-likelihood, perplexity, and generation. The log-likelihood is more suitable for multiple-choice tasks since it helps to restrict the model’s option to fewer choices—more appropriate for weaker models. However, the log-likelihood approach cannot evaluate the generative capabilities of LLMs to generate coherent and relevant answers. Moreover, closed models are only accessible via API and do not provide access to the log probabilities, making it impossible to use the log-likelihood approach. To extract the correct answers for the task, we employed a verbalizer [25, 52]. For AfriMGSM, we used the default verbalizer provided by the tool. However, for AfriXNLI and AfriMMLU, we manually created a verbalizer for the closed models and used the log-likelihood request type for the open models.Table 8 provides the prompt templates used for evaluation present in the tool. For XNLI, we employed two types of prompts to understand the sensitivity of evaluation to type of prompt. We report on these results in Table 6.
Table 8: Prompt templates used for different tasks and datasets
. Task/Dataset Prompt AfriMGSM Question: {question} Answer: AfriMMLU You are a highly knowledgeable and intelligent artificial intelligence model answers multiple-choice questions about {subject} Question: {question} Choices: A: {choice1} B: {choice2} C: {choice3} D: {choice4} Answer: AfriXNLI (ANLI prompt [44]) {premise} Question: {hypothesis} True, False, or Neither? Answer: AfriXNLI (Lai’s prompt [37]) Please identify whether the premise entails or contradicts the hypothesis in the following premise and hypothesis. The answer should be exact entailment, contradiction, or neutral. Premise: {premise} Hypothesis: {hypothesis} Is it entailment, contradiction, or neutral?
@@ A.3 AfriCOMET metric scores for XNLI translation
We employ AfriCOMET evaluation metrics, as developed by Wang et al. [62], to automatically assess the quality of translations for our newly created benchmarks. Figure 3 depicts the histogram of scores obtained from AfriCOMET for AfriXNLI, illustrating promising results and offering compelling evidence for the effectiveness of our translations (Amharic, Yorùbá, isiZulu). However, the performance of this metric depends on if the language we are evaluating is covered in the pre-training of the base model of the metric i.e. AfroXLMR-large. In the case of Lingala, Twi and Wolof, the performance of the metric does not correlate with the human translation since they are not covered in AfroXLMR. Similar findings were reported in the original AfriCOMET QE paper that Twi had worse correlation with human judgement (i.e., 0.279 for Pearson, and 0.060 for Spearman) [62].

### v2

@@ 4.1 Overall Results
Figure 2: Few-shot evaluation on IrokoBench, we performed 8-shots for AfriMGSM and 5-shots for the others.
AfriXNLI AfriMMLU AfriMGSM
Model t1 t2 t3 t4 t5 Ave. t1 t2 t3 t4 t5 Ave. t1 t2 t3 t4 t5 Ave.
Aya-101 47.7 49.9 45.7 51.5 48.3 ${48.6}_{\pm 2.2}$ 29.6 29.7 29.7 29.7 29.6 ${29.7}_{\pm 0.1}$ 4.4 4.1 4.2 4.4 4.3 ${4.3}_{\pm 0.1}$
Gemma 2 27B 38.3 40.3 33.4 34.1 35.4 ${36.3}_{\pm 2.9}$ 39.9 39.2 38.7 39.2 39.1 ${39.2}_{\pm 0.4}$ 22.3 26.3 25.0 27.0 26.2 ${25.4}_{\pm 1.9}$
GPT-4o 55.3 49.6 64.3 62.9 59.7 ${58.4}_{\pm 6.0}$ 60.0 55.8 59.0 39.6 59.6 ${54.8}_{\pm 8.7}$ 46.7 49.8 48.0 49.5 49.6 ${48.7}_{\pm 1.3}$
Table 7: Ablation results: Effect of using different prompts (t1, t2, t3, t4, and t5) for IrokoBench datasets. Best prompt results are in bold. Average computed on only African languages. Second best prompt are underlined.
4.2 Task-specific results
Here, we examine the individual language performance per task, comparing which task prefers in-language v.s. translate-test evaluation. We compare the performance on a subset of LLMs: Aya-101, Gemma 2 27B LLaMa 3.1 70B, and GPT-4o. Other LLMs are in Appendix A.4.
AfriMMLU evaluation is better when prompting in-language for closed LLMs
Table 3 shows the result of different LLMs on AfriMMLU using in-language vs. translate-test. For GPT-4o, we found in-language prompting to be generally better. However, for open models like LLaMa 3.1 70B and Gemma 2 27B, we find a large improvement in the performance of the translate-test for 15 out of 16 African languages. The only language that did not improve is wol, probably due to poor MT performance on NLLB-200, as reported in Costa-jussà et al. (2024). This shows the benefit of translate-test for prompting English-centric LLMs when evaluating low-resource languages.
AfriMMLU performance by subjects
Table 4 shows the result of GPT-4o by subjects. Interestingly, elementary math achieved the best overall accuracy, where 13 out of 16 African languages achieved at least 70% despite struggling with AfriMGSM. The performance difference to AfriMGSM may be due to the multi-choice options of MMLU, which may be slightly easier than free-form answer. International law also achieved impressive performance with 10 out of 16 languages achieving 70%. All languages struggle the most with Global facts including English and French. Similarly, many African languages find it difficult to answer questions in geography and microeconomics subjects. This presents an opportunity for improving LLMs for the education domain.
AfriMGSM performance receives significant boost with translate-test for open models
Table 5 shows that we can achieve a significant boost in performance with the translate-test on all languages we evaluated. We hypothesize that the current LLMs are better at reasoning in English than other languages. Gemma 2 27B and LLaMa 3.1 70B improved by $+17.6$ and $+21.0$ respectively when questions are asked in English, while closed models like GPT-4o dropped in performance.
Cross-lingual transfer in-language achieves better results
When there is large enough labeled data in English, we could leverage this cross-lingual signal for zero-shot evaluation. We trained on 400k English NLI examples, and we performed zero-shot transfer in in-language and translate-test setting. Table 6 shows that cross-lingual transfer using an Africa-centric smaller language model (AfroXLMR-76L) gave better results than prompting LLMs on average. AfroXLMR-76L has been pre-trained on all languages in AfriXNLI, which explains the impressive performance. However, for multilingual encoders that have not seen some of the languages, prompting GPT-4o seems to be better, as shown in Appendix A.6.
Impact of few-shot vs zero-shot
Figure 2 shows the few-shot results for the IrokoBench datasets leveraging Gemma 2 27B, LLaMa 3.1 70B and GPT-4o when we provide few examples in in-language setting. We found out that Gemma 2 27B and LLaMa 3.1 70B LLM consistently benefited the least from additional few shots examples for classification tasks (AfriXNLI and AfriMMLU) where Gemma 2 27B improved by $+13.2$ and $+4.9$ respectively. Similarly, LLaMa 3.1 70B improved on both AfriXNLI and AfriMMLU tasks with $+10.7$ and $+7.0$. However, for reasoning tasks, only GPT-4o improved in performance by $6.3$, other LLMs dropped in performance, probably due to their inability to reason in non-English languages. Surprisingly, GPT-4o did not benefit from additional examples for the classification tasks.
Prompt t2
{{premise}}
Question: {hypothesis} True, False, or Neither?
Answer:
Prompt t3
Given the following premise and hypothesis in English, identify if the premise entails, contradicts, or is neutral towards the hypothesis. Please respond with exact ’entailment’, ’contradiction’, or ’neutral’.
Premise: {premise}
Hypothesis: {hypothesis}
Prompt t4
You are an expert in Natural Language Inference (NLI) specializing in the {Language} language. Analyze the premise and hypothesis given in {Language}, and determine the relationship between them. Respond with one of the following options: ‘’entailment’, ’contradiction’, or ’neutral’.
Premise: {premise}
Hypothesis: {hypothesis}
Table 8: Three different prompts preferred by different models for AfriXNLI
Sensitivity to prompt templates
To understand whether sensitivity to prompts impacts results, we perform an ablation for all the IrokoBench tasks where we evaluate the performance of five different prompts (see subsection A.2). Table 7 shows the results of five prompts we tested for (three of the prompts most preferred by different models are shown in Table 8). On AfriXNLI, we find that simpler prompt (Nie et al., 2020) i.e. {{premise}} Question: {{hypothesis}} True, False, or Neither?, have better results for the open models like Aya-101 and Gemma 2 27B, while GPT-4o prefers t3 where a detailed task description is provided. The best prompt for Aya-101 is t4 where the {{language}} name is mentioned, which shows additional language information may be useful in improving performance.
On AfriMMLU, we find GPT-4o perform worse for t4.444Analyze each question critically and determine the most correct option based on your understanding of the subject matter ” Question: {question}. Choices A: {choice1}, B:{choice2}, C: {choice3}, D: {choice4} . However, other models are not very sensitive to the use of different prompts. In general, we do not find AfriMGSM to be sensitive to different prompts. In subsection A.4, we provide the results of five prompt templates for all LLMs evaluated.
@@ 5 Conclusion
Our benchmark has a few limitations: (1) The benchmark is human-translated which may include some translationese effects, it would have been better if they are all generated in the native African languages. However, this parallel translation allows us to evaluate and compare the same sentences in all these languages. (2) We only cover three language families in Africa, Nilo-Saharan, Austronesian, and Khoisan language groups are missing, one of the reason we excluded them is either lack of contact with professional translators or limited translation budget, we hope to extend to more languages in the future.
@@ Acknowledgment
This work was supported by Lacuna Fund, an initiative co-founded by The Rockefeller Foundation, Google.org, and Canada’s International Development Research Centre. Additional support was provided through compute credits from Oracle and the Cohere For AI Research Grant. We extend our gratitude to Charles Riley for facilitating our connection with the Vai translator for the MGSM data translation.
We are also deeply thankful to OpenAI for granting API credits through their Researcher Access API program to Masakhane, enabling the evaluation of GPT-3.5 and GPT-4 LLMs. Similarly, we appreciate Google for providing GCP credits via the Gemma 2 Academic Program, which supported the Gemini-1.5-Pro inference. Lastly, we would like to thank Hailey Schoelkopf and Lintang Sutawika for their invaluable assistance with the EleutherAI lm-eval tool.
Appendix A Appendix
@@ A.1 Language covered
Table 9 provide the languages covered in the IrokoBench, their language family, regions located in Africa, number of speakers, and size of monolingual data available on the web based on the MADLAD cleaned corpus (Kudugunta et al., 2023)—we only report number of characters in mega bytes. Additionally, we added an indication whether this language is covered in the pre-training of Aya-101 and BLOOMZ 7B LLMs.
A.2 Prompts Template and Evaluation Tool
@@
We use the EleutherAI LM Evaluation Harness (lm-eval) tool (Biderman et al., 2024)—a popular evaluation tool that is helping to standardize LLM evaluation. The tool allows for three types of evaluation: log-likelihood, perplexity, and generation. The log-likelihood is more suitable for multiple-choice tasks since it helps to restrict the model’s option to fewer choices—more appropriate for weaker models. However, the log-likelihood approach cannot evaluate the generative capabilities of LLMs to generate coherent and relevant answers. Moreover, closed models are only accessible via API and do not provide access to the log probabilities, making it impossible to use the log-likelihood approach. To extract the correct answers for the task, we employed a verbalizer Gao et al. (2021); Schick and Schütze (2021). For AfriMGSM, we used the default verbalizer provided by the tool. However, for AfriXNLI and AfriMMLU, we manually created a verbalizer for the closed models and used the log-likelihood request type for the open models.
The prompt templates used for evaluation of different tasks are in Table 11, Table 12 and Table 10.
Prompt1
Question: {question}
Answer:
Prompt2
Give direct numerical answers for the question provided.
Question: {question}
Answer:
Prompt3
Solve the following math question.
Question: {question}
Answer:
Prompt4
Answer the given question with the appropriate numerical value, ensuring that the response is clear and without any supplementary information.
Question: {question}
Answer:
Prompt5
For mathematical questions provided in language language. Supply the accurate numeric answer to the provided question
Question: {question}
Answer:
Table 10: Five different prompt used for prompt sensitivity experiments in AfriMGSM
Prompt 1
Please identify whether the premise entails or contradicts the hypothesis in the following premise and hypothesis. The answer should be exact entailment, contradiction, or neutral.
Premise: {premise}
Hypothesis: {hypothesis}
Is it entailment, contradiction, or neutral?
Prompt 2
{{premise}}
Question: {hypothesis} True, False, or Neither?
Answer:
Prompt 3
Given the following premise and hypothesis in English, identify if the premise entails, contradicts, or is neutral towards the hypothesis. Please respond with exact ’entailment’, ’contradiction’, or ’neutral’.
Premise: {premise}
Hypothesis: {hypothesis}
Prompt 4
You are an expert in Natural Language Inference (NLI) specializing in the {Language} language. Analyze the premise and hypothesis given in {Language}, and determine the relationship between them. Respond with one of the following options: ‘’entailment’, ’contradiction’, or ’neutral’.
Premise: {premise}
Hypothesis: {hypothesis}
Prompt 5
Based on the given statement, is the following claim ’true’, ’false’, or ’inconclusive’.
Premise: {premise}
Hypothesis: {hypothesis}
Table 11: Five different prompt used for prompt sensitivity experiments in AfriXNLI
Prompt 1
You are a highly knowledgeable and intelligent artificial intelligence model answers multiple-choice questions about {subject}
Question: {question}
A: {choice1}
B: {choice2}
C: {choice3}
D: {choice4}
Answer:
Prompt 2
As an expert in {subject}, choose the most accurate answer to the question below. Your goal is to select the correct option ’A’, ’B’, ’C’, or ’D’ by understanding the nuances of the topic.
Question: {question}
A: {choice1}
B: {choice2}
C: {choice3}
D: {choice4}
Answer:
Prompt 3
You are a subject matter expert in {subject}. Utilizing your expertise in {subject}, answer the following multiple-choice question by picking ’A’, ’B’, ’C’, or ’D’.
Question: {question}
A: {choice1}
B: {choice2}
C: {choice3}
D: {choice4}
Answer:
Prompt 4
Analyze each question critically and determine the most correct option based on your understanding of the subject matter
A: {choice1}
B: {choice2}
C: {choice3}
D: {choice4}
Answer:
Prompt 5
Given your proficiency in {subject}, please answer the subsequent multiple-choice question with ’A’, ’B’, ’C’, or ’D’.
Question: {question}
A: {choice1}
B: {choice2}
C: {choice3}
D: {choice4}
Answer:
Table 12: Five different prompt used for prompt sensitivity experiments in AfriMMLU
@@ A.3 AfriCOMET metric scores for XNLI translation
We employ AfriCOMET evaluation metrics, as developed by Wang et al. (2024), to automatically assess the quality of translations for our newly created benchmarks. Figure 3 depicts the histogram of scores obtained from AfriCOMET for AfriXNLI, illustrating promising results and offering compelling evidence for the effectiveness of our translations (Amharic, Yorùbá, isiZulu). However, the performance of this metric depends on if the language we are evaluating is covered in the pre-training of the base model of the metric i.e. AfroXLMR-large. In the case of Lingala, Twi and Wolof, the performance of the metric does not correlate with the human translation since they are not covered in AfroXLMR. Similar findings were reported in the original AfriCOMET QE paper that Twi had worse correlation with human judgement (i.e., 0.279 for Pearson, and 0.060 for Spearman) (Wang et al., 2024).